# Bitangent surfaces and involutions of quartic surfaces

### v3

@@ 1. Introduction
Let $G_{1}(\mathbb{P}^{3})$ be the Grassmannian of lines in $\mathbb{P}^{3}$. For a point in $G_{1}(\mathbb{P}^{3})$, the corresponding line in $\mathbb{P}^{3}$ is called a ray. A surface $F\subset G_{1}(\mathbb{P}^{3})$ is called a congruence of lines. Its algebraic class $[F]$ in the Chow ring $A^{\bullet}(G_{1}(\mathbb{P}^{3}))$ is determined by two numbers: the order $m$ and the class $n$. The order (resp. class) is the number of lines on $F$ passing through a general point (resp. contained in a general plane) in $\mathbb{P}^{3}$. We will call the pair $(m,n)$ the bidegree of $F$. The number $m+n$ is equal to the degree of $F$ in the Plücker embedding $G_{1}(\mathbb{P}^{3})\hookrightarrow\mathbb{P}^{5}$ (cf. [20, Chapter X, §2], [3], [11]).
@@
For example, it is known that, for a general surface of degree $d$, the bidegree of the surface ${\mathrm{Bit}}(X)$ is equal to $(\frac{1}{2}d(d-1)(d-2)(d-3),\frac{1}{2}d(d-2)(d^{2}-9))$ [19, Volume 1, p. 281]. In particular, the bidegree of the bitangent surface of a general quartic surface is equal to $(12,28).$ Moreover, the surface is a smooth surface of degree $40$ in $\mathbb{P}^{5}$. It is of general type with $p_{g}=45$ and $K^{2}=360$ [24], [27].
@@
Although admitting some mild singular points in $X$ doesn’t change the bidegree, it may drastically change the surface ${\mathrm{Bit}}(X)$, it even may become reducible. For example, when $X$ is realized as the focal surface of a congruence of lines of bidegree $(2,n)$ (the focal surface of a congruence $F$ of lines of order $n>1$ is a surface in $\mathbb{P}^{3}$ such that all rays of $F$ are its bitangents, see, for example, [20, Chapter X, §2], the surface ${\mathrm{Bit}}(X)$ is always singular although its singular points are ordinary double points. In the case $n=2$, the congruence of lines is a del Pezzo surface of degree $4$ and the focal surface is the famous $16$-nodal Kummer quartic surface $X$ [15], and the number of irreducible components of ${\mathrm{Bit}}(X)$ is equal to 22, six of them are of bidegree $(2,2)$ and 16 of them are of bidegree $(0,1)$. They are planes of lines contained in one of the 16 planes intersecting $X$ along a double conic, a trope plane.
@@ 2. Generalities on the surface of bitangent lines
Let ${\mathrm{Bit}}(X)\subset G_{1}(\mathbb{P}^{3})$ be the variety of bitangent lines.
@@
An irreducible surface $S$ in $G_{1}(\mathbb{P}^{3})$ of bidegree $(m,0)$ (resp. $(0,n)$) is an $\alpha$-plane (resp. $\beta$-plane).
@@
Passing to the dual congruence of lines in the dual space of $\mathbb{P}^{3}$, it is enough to prove that any irreducible congruence of bidegree $(0,n)$ is an $\beta$-plane. Let $Z_{S}=\{(x,\ell)\in\mathbb{P}^{3}\times G_{1}(\mathbb{P}^{3}):x\in\ell\}$ be the restriction of the tautological projective bundle over $G_{1}(\mathbb{P}^{3})$ to $S$. Since $S$ is irreducible, $Z_{S}$ is an irreducible $3$-fold. Since $m=0$, its image $Y$ under the first projection is an irreducible subvariety of dimension $\leq 2$, the union of rays of $S$ that pass through some point in $Y$. Suppose $\dim Y=1$. Then, the fibers of $Z_{S}\to Y$ are $\alpha$-planes that sweep at least one plane. So, $Y$ cannot be a curve. Thus, $Y$ is a surface. The image of the fiber of $Z_{S}\to S$ over a point $y\in Y$ in $S$ is a curve contained in the $\alpha$-plane $\Omega(y)$ of lines through $y$. The corresponding rays sweep a cone with vertex at $y$ contained in $Y$. Thus, $Y$ is a cone at each of its points. Since $Y$ is reduced, this can happen only if $Y$ is a plane. So, all rays of $S$ are contained in a plane, and hence, $S$ is an $\beta$-plane. ∎
@@
$1\leq m\leq\frac{1}{2}d(d-1)(d-2)(d-3),\quad 1\leq n=\frac{1}{2}d(d-2)(d^{2}-9).$
@@
We follow Salmon [19, Volume 2, p. 281]. Let $q$ be a general point in $\mathbb{P}^{3}$ and let $\ell=\langle q,q^{\prime}\rangle$ be the line containing $q$ and tangent to $X$ at some point $q^{\prime}=[x_{0},y_{0},z_{0},w_{0}]$. Without loss of generality, we may assume that $q=[0,0,0,1]$ and
@@
The fiber of the first projection consists of bitangent lines $\ell$ passing through $q$. By above, it consists of finitely many lines. This implies that each irreducible component of $M$ is of dimension $3$. Since the fibers of the second projections are lines, the image of each irreducible component of $M$ is a surface. ∎
We denote by ${\mathrm{Flex}}(X)\subset G_{1}(\mathbb{P}^{3})$ the variety of lines in $\mathbb{P}^{3}$ that either contained in $X$ or intersect $X$ at some point with multiplicity $\geq 3$. It intersects ${\mathrm{Bit}}(X)$ at the set of lines contained in $X$ and the set of bitangents that intersect $X$ at one point.
@@
We use the following theorem due to Valentine and Viktor Kulikov [14, Theorem 0.1]:
@@ 3. Quartic surfaces: $p\neq 2$
It follows from Kummer’s classification of congruences of order $2$ without fundamental curves that the moduli space of quartic surfaces with $16\geq\mu\geq 11$ nodes contains an irreducible component such that a general surface from this component has reducible congruence ${\mathrm{Bit}}(X)$. These irreducible components are realized as the focal surfaces of the congruences of order 2 without fundamental curves (for example, as we will explain below, the first case in the following Table 2 is a quartic del Pezzo surface). The number of irreducible components and their bidegree are given in Table 2 below (see [15], [23, Art. 348]).
@@
We have already mentioned in the Introduction that a Kummer surface contains 16 planes cut out the surface along a smooth conic. Recall that the Kummer quartic surface is the quotient of the Jacobian of a smooth curve of genus two by the inversion automorphism. The sixteen double conics (tropes) are the images of the theta divisor and its translations by sixteen 2-torsion points. This is a classical, well-known fact that can be found in any treatment of Kummer surfaces (see, for example, [3, 10.3], [9]). Let us explain how to see the six irreducible components of bidegree $(2,2)$. Recall that $X$ admits a smooth model $\tilde{X}$ as a surface of degree $8$ in $\mathbb{P}^{5}$ (see [3, 10.3.3]). The surface $\tilde{X}$ is a complete intersection of three quadrics in $\mathbb{P}^{5}$ which is a $K3$ surface. One can choose projective coordinates such that $\tilde{X}$ is given by the following equations:
@@
The first quadric $V(\sum z_{i}^{2})$ can be identified with the Grassmannian quadric $G_{1}(\mathbb{P}^{3})$ written in Klein coordinates $z_{i}$ corresponding to six apolar line complexes (see, for example, [3, 10.2]). It is known that any automorphism of $G_{1}(\mathbb{P}^{3})$ comes from a projective collineation or correlation. The involutions $g_{i}$ preserve the apolar line complexes $V(z_{i})$, and hence, come from a correlation. They transform points to planes, and therefore, act on $G_{1}(\mathbb{P}^{3})$ by transforming the plane of lines through a point (an $\alpha$-plane) to the plane of lines contained in a given plane (a $\beta$-plane).
@@
First, we recall the relation between $X,\tilde{X}$ with the quadratic line complex $Y=V(\sum a_{i}z_{i}^{2})\cap G_{1}(\mathbb{P}^{3})$ (see, e.g., [3, §10.3.3]). For a point $x\in\mathbb{P}^{3}$, let $\Omega(x)\subset G_{1}(\mathbb{P}^{3})$ be the $\alpha$-plane consisting of lines passing through $x$, and for a plane $\Pi\subset\mathbb{P}^{3}$, denote by $\Omega(\Pi)$ the $\beta$-plane consisting of lines lying on $\Pi$. The Kummer quartic surface $X$ is the set of $x\in\mathbb{P}^{3}$ such that $\Omega(x)\cap Y$ is a singular conic in $\Omega(x)$. The surface $\tilde{X}$ sits in $G_{1}(\mathbb{P}^{3})$ as a singular surface of the quadratic line complex, that is, it parameterizes the lines that are singular points of the intersection of $Y$ with $\alpha$-planes. We can identify each point $x\in X$ with the $\alpha$-plane $\Omega(x)$. An automorphism $g_{i}$ acts on $X$ by sending $\Omega(x)$ to $\Omega(\Pi)$. Let $x\in X$ be a general point and let $\Pi=\mathbb{T}_{x}(X)$ be the embedded tangent space of $X$ at $x$. As we explained above, $g_{i}$ interchanges the two families $\{\Omega(x)\}_{x\in\mathbb{P}^{3}}$, $\{\Omega(\Pi)\}_{\Pi\subset\mathbb{P}^{3}}$ of planes, and hence,
@@ 4. Birational involutions of a quartic surface
where $P=[a,b,c,d],x=[x_{0},x_{1},x_{2},x_{3}],T(x)=[y_{0},y_{1},y_{2},y_{3}]$. It is expected to be a curve of degree $12k+$ 1, where $k$ is the algebraic degree of $T$ [10, p. 175]. However, it also includes the locus of fundamental points of $T$, as well as the closure of the locus of fixed points of $T$. The class of $S(\sigma)$ is expected to be equal to the degree of $\Pi\cap T(\Pi)\cap X$ for a general plane $\Pi$, which is $4k$.
@@ 5. Quartic surfaces: $p=2$
First of all, ${\mathrm{Bit}}(X)$ may not be a surface. For example, let $X$ be a quartic surface in characteristic two with the equation:
$F=x_{0}^{4}+x_{0}^{2}(x_{1}^{2}+A_{2}(x_{2},x_{3}))+x_{1}^{4}+B_{4}(x_{2},x_{3})=0,$
where $A_{2},B_{4}$ are homogeneous polynomials in $x_{2},x_{3}$ of degree 2 and 4, respectively. Let $x=[a,b,0,0]\in V(x_{2},x_{3})$. Substituting in $F$ the parametric equation of a line passing through $x$, we find that the projection map with the center at $x$ is inseparable. Thus, ${\mathrm{Bit}}(X)$ contains a line of $\alpha$-planes.
A point $x$ in $\mathbb{P}^{3}$ is called an inseparable projection center of a normal surface $X$ if the projection map with the center at $x$ is inseparable.
@@
The set of inseparable projection centers of a normal quartic surface in characteristic $2$ is a finite set or consists of a line. Any inseparable projection center contained in $X$ is a singular point of $X$.
@@
Since the set of points $x\in X$ such that the line $\langle q,x\rangle$ is tangent at $x$ is equal to the intersection $X\cap P_{q}(X)$ [3, Theorem 1.1.5], a point $q=[a_{0},a_{1},a_{2},a_{3}]$ is an inseparable center of a quartic surface $X=V(F)$ if and only if the polar $P_{q}(X)=V(\sum a_{i}\frac{\partial F}{\partial x_{i}})=\mathbb{P}^{3}$. Suppose the set of inseparable projection centers contains curve of degree $\geq 2$ (not necessary irreducible). Then we can choose three non-collinear points on it. Choose projective coordinates such that these points are $[1,0,0,0],[0,1,0,0]$ and $[0,0,1,0]$. Then, the partials $\frac{\partial F}{\partial x_{i}}=0$ for $i=0,1,2$. This means that $F$ contains $x_{0},x_{1},x_{2}$ in even power. But, then, it must contain $x_{3}$ in even power. Hence, $F$ is a square.
@@
Suppose $X$ is nonsingular, then each irreducible componen of ${\mathrm{Bit}}(X)$ is a surface.
@@
By Corollary 2.5, any surface in characteristic zero with ordinary double points as singularities is a general projection surface. In the next section, we show that it is not true anymore in characteristic $2$.
@@
where $Q$ is a quadratic form and $F$ is a cubic form. The singular locus of $X$ is equal to $V(L,Q,F)\cup V(Q,\textrm{Sing}(V(F))$. In particular, $X$ is a singular quartic surface.
@@
Without loss of generality, we may assume that $L=x$ is a coordinate plane. Taking the partials, we find the singular locus is indeed $V(x,Q,F)$. ∎
Recall that the $p$-rank of a smooth curve $C$ of genus $g>0$ over an algebraically closed field of characteristic $p>0$ is the $p$-rank of the elementary abelian group $J(C)[p]$ of $p$-torsion points of its jacobian variety of $C$. It takes values in the set $[0,g]$. When $p=2$ and $C$ is a smooth plane quartic, the $p$-rank takes values in $\{3,2,1,0\}$ and can be characterized by the number of bitangents of $C$ equal to $7,4,2,1$, respectively [18], [22, §3].
The proof of the following proposition can be found in [26, Proposition 1] (see also [18]).
@@
Computing the stabilizer group of projective transformations that leave the normal form (1)–(4) invariant, we see that smooth quartics of $2$-rank less than $3$ form three hypersurfaces in the space of all smooth plane quartics.
It is clear that there exist smooth quartic surfaces with a general plane section isomorphic to a plane quartic curve with $p$-rank equal to $3$, we do not know whether it is true for plane curves of rank 2 and 1. The next proposition shows that it is not true for supersingular plane quartic curves.
@@
Let $X$ be a normal quartic surface whose general hyperplane section is isomorphic to a plane quartic with the $2$-rank equal to 00. Then, ${\mathrm{Bit}}(X)$ contains a $\beta$-plane, and hence, it is singular.
@@
$F:=\{(\ell,\Pi)\in{\mathrm{Bit}}(X)\times\check{\mathbb{P}}^{3}:\ell\subset\Pi\},$
and let $f:F\to\check{\mathbb{P}}^{3}$ be the second projection. Consider a general plane section as a plane curve over a field of rational functions of the dual projective space ˇ$\check{\mathbb{P}}^{3}$. Applying Proposition 5.8, we find a family $f^{\prime}:F^{\prime}\to B$ where $B$ is a finite cover of ˇ$\check{\mathbb{P}}^{3}$ such that $F^{\prime}$ contains an irreducible component that is mapped to a point. This corresponds to the plane $V(x)$ in equation (4) from Proposition 5.8. The image of this component in $F$ is a $\beta$-plane $\Pi$ contained in ${\mathrm{Bit}}(X)$. Further computation shows that there is nothing else. ∎
@@
The surface $X$ has a unique (non-rational) singular point $[0,0,0,1]$ and an ordinary node $[0,0,1,0]$. The point $[0,0,0,1]$ is a unique inseparable projection center of $X$. In fact, take a general point $P=[x_{0},y_{0},z_{0},w_{0}]$ in $\mathbb{P}^{3}$. Substituting the parametric equation $[x,y,z,w]=[sx_{0}+tu_{0},sy_{0}+tu_{1},sz_{0}+tu_{2},sw_{0}+tu_{3}]$, we find one bitangent line that connects $P$ with the point $[0,0,0,1]$. Thus, ${\mathrm{Bit}}(X)$ contains an $\alpha$-plane. On the other hand, a general hyperplane section $w=\alpha x+\beta y+\gamma z$ is the case of Proposition 5.8, that is, it is a plane quartic curve with the 2-rank 0. By Proposition 5.9, it contains a $\beta$-plane. Thus, ${\mathrm{Bit}}(X)$ contains the union of an $\alpha$-plane and a $\beta$-plane. Further computation shows that there is nothing else.

### v4

@@ 1. Introduction
Let $G_{1}(\mathbb{P}^{3})$ be the Grassmannian of lines in $\mathbb{P}^{3}$. For a point in $G_{1}(\mathbb{P}^{3})$, the corresponding line in $\mathbb{P}^{3}$ is called a ray. A surface $F\subset G_{1}(\mathbb{P}^{3})$ is called a congruence of lines. Its algebraic class $[F]$ in the Chow ring $A^{\bullet}(G_{1}(\mathbb{P}^{3}))$ is determined by two numbers: the order $m$ and the class $n$. The order (resp. class) is the number of lines on $F$ passing through a general point (resp. contained in a general plane) in $\mathbb{P}^{3}$. We will call the pair $(m,n)$ the bidegree of $F$. The number $m+n$ is equal to the degree of $F$ in the Plücker embedding $G_{1}(\mathbb{P}^{3})\hookrightarrow\mathbb{P}^{5}$ (cf. [22, Chapter X, §2], [3], [12]).
@@
For example, it is known that, for a general surface of degree $d$, the bidegree of the surface ${\mathrm{Bit}}(X)$ is equal to $(\frac{1}{2}d(d-1)(d-2)(d-3),\frac{1}{2}d(d-2)(d^{2}-9))$ [21, Volume 1, p. 281]. In particular, the bidegree of the bitangent surface of a general quartic surface is equal to $(12,28).$ Moreover, the surface is a smooth surface of degree $40$ in $\mathbb{P}^{5}$. It is of general type with $p_{g}=45$ and $K^{2}=360$ [26], [29].
@@
Although admitting some mild singular points in $X$ doesn’t change the bidegree, it may drastically change the surface ${\mathrm{Bit}}(X)$, it even may become reducible. For example, when $X$ is realized as the focal surface of a congruence of lines of bidegree $(2,n)$ (the focal surface of a congruence $F$ of lines of order $n>1$ is a surface in $\mathbb{P}^{3}$ such that all rays of $F$ are its bitangents, see, for example, [22, Chapter X, §2]), the surface ${\mathrm{Bit}}(X)$ is always singular although its singular points are ordinary double points. In the case $n=2$, the congruence of lines is a del Pezzo surface of degree $4$ and the focal surface is the famous $16$-nodal Kummer quartic surface $X$ [16], and the number of irreducible components of ${\mathrm{Bit}}(X)$ is equal to 22, six of them are of bidegree $(2,2)$ and 16 of them are of bidegree $(0,1)$. They are planes of lines contained in one of the 16 planes intersecting $X$ along a double conic, a trope plane.
@@ 2. Generalities on the surface of bitangent lines
Recall that $G_{1}(\mathbb{P}^{3})$ contains two types of planes, one is called an $\alpha$-plane consisting of lines passing through a point of $\mathbb{P}^{3}$ and other is called a $\beta$-plane consisting of lines contained in a plane of $\mathbb{P}^{3}$.
@@
An irreducible integral surface $S$ in $G_{1}(\mathbb{P}^{3})$ of bidegree $(m,0)$ (resp. $(0,n)$) is an $\alpha$-plane (resp. $\beta$-plane). In particular, $m=1$ (resp. $n=1$).
@@
Passing to the dual congruence of lines in the dual space of $\mathbb{P}^{3}$, it is enough to prove that any irreducible congruence of bidegree $(0,n)$ is a $\beta$-plane. Let $Z_{S}=\{(x,\ell)\in\mathbb{P}^{3}\times S:x\in\ell\}$ be the restriction of the tautological projective bundle over $G_{1}(\mathbb{P}^{3})$ to $S$. Since $S$ is irreducible, $Z_{S}$ is an irreducible $3$-fold. Since $m=0$, its image $Y$ under the first projection is an irreducible subvariety of dimension $1$ or $2$. Suppose $\dim Y=1$. Then, the fibers of $Z_{S}\to Y$ are $\alpha$-planes that sweep at least one plane, and hence $S$ contains an $\alpha$-plane. So, $Y$ cannot be a curve. Thus, $Y$ is a surface. The image of the fiber of $Z_{S}\to S$ over a point $y\in Y$ in $S$ is a curve contained in the $\alpha$-plane $\Omega(y)$ of lines through $y$. The corresponding rays sweep a cone with vertex at $y$ contained in $Y$. Thus, $Y$ is a cone at each of its points. Since $Y$ is reduced, this can happen only if $Y$ is a plane. So, all rays of $S$ are contained in a plane, and hence, $S$ is a $\beta$-plane. ∎
@@
$1\leq m\leq\frac{1}{2}d(d-1)(d-2)(d-3),\quad n=\frac{1}{2}d(d-2)(d^{2}-9).$
@@
We follow Salmon [21, Volume 2, p. 281]. Let $q$ be a general point in $\mathbb{P}^{3}$ and let $\ell=\langle q,q^{\prime}\rangle$ be the line containing $q$ and tangent to $X$ at some point $q^{\prime}=[x_{0},y_{0},z_{0},w_{0}]$. Without loss of generality, we may assume that $q=[0,0,0,1]$ and
@@
The fiber of the first projection consists of bitangent lines $\ell$ passing through $q$. By the above, it consists of finitely many lines. This implies that each irreducible component of $M$ is of dimension $3$. Since the fibers of the second projections are lines, the image of each irreducible component of $M$ is a surface. ∎
We denote by ${\mathrm{Flex}}(X)\subset G_{1}(\mathbb{P}^{3})$ the variety of lines in $\mathbb{P}^{3}$ that are either contained in $X$ or intersect $X$ at some point with multiplicity $\geq 3$. It intersects ${\mathrm{Bit}}(X)$ at the set of lines contained in $X$ and the set of bitangents that intersect $X$ at one point.
@@
We use the following theorem due to Valentine and Viktor Kulikov [15, Theorem 0.1]:
@@ 3. Quartic surfaces: $p\neq 2$
It follows from Kummer’s classification of congruences of order $2$ without fundamental curves that the moduli space of quartic surfaces with $16\geq\mu\geq 11$ nodes contains an irreducible component such that a general surface from this component has reducible congruence ${\mathrm{Bit}}(X)$. Some of the irreducible components are confocal congruences, i.e., congruences of the same bidegree that share the same focal surface (for example, as we will explain below, the first case in the following Table 2 is a quartic del Pezzo surface). The number of irreducible components and their bidegree are given in Table 2 below (see [16], [25, Art. 348]).
@@
We have already mentioned in the introduction that a Kummer surface admits 16 planes cutting the surface along a smooth conic. Recall that a Kummer quartic surface is the quotient of the Jacobian of a smooth curve of genus two by the inversion automorphism. The sixteen double conics (tropes) are the images of the theta divisor and its translations by sixteen 2-torsion points. This is a classical, well-known fact that can be found in any treatment of Kummer surfaces (see, for example, [3, 10.3], [10]). Let us explain how to see the six irreducible components of bidegree $(2,2)$. Recall that $X$ admits a smooth model $\tilde{X}$ as a surface of degree $8$ in $\mathbb{P}^{5}$ (see [3, 10.3.3]). The surface $\tilde{X}$ is a complete intersection of three quadrics in $\mathbb{P}^{5}$ which is a $K3$ surface. One can choose projective coordinates such that $\tilde{X}$ is given by the following equations:
@@
The first quadric $V(\sum z_{i}^{2})$ can be identified with the Grassmannian quadric $G_{1}(\mathbb{P}^{3})$ written in Klein coordinates $z_{i}$ corresponding to six apolar line complexes (see, for example, [3, 10.2]). It is known that any automorphism of $G_{1}(\mathbb{P}^{3})$ comes from a projective collineation or correlation. The involutions $g_{i}$ preserve the apolar line complexes $V(z_{i})$, and, hence, come from a correlation. They transform points to planes, and therefore, act on $G_{1}(\mathbb{P}^{3})$ by transforming the plane of lines through a point (an $\alpha$-plane) to the plane of lines contained in a given plane (a $\beta$-plane).
@@
First, we recall the relation between $X,\tilde{X}$ and the quadratic line complex $Y=V(\sum a_{i}z_{i}^{2})\cap G_{1}(\mathbb{P}^{3})$ (see, e.g., [3, §10.3.3]). For a point $x\in\mathbb{P}^{3}$, let $\Omega(x)\subset G_{1}(\mathbb{P}^{3})$ be the $\alpha$-plane consisting of lines passing through $x$, and for a plane $\Pi\subset\mathbb{P}^{3}$, denote by $\Omega(\Pi)$ the $\beta$-plane consisting of lines lying on $\Pi$. The Kummer quartic surface $X$ is the set of $x\in\mathbb{P}^{3}$ such that $\Omega(x)\cap Y$ is a singular conic in $\Omega(x)$. The surface $\tilde{X}$ sits in $G_{1}(\mathbb{P}^{3})$ as a singular surface of the quadratic line complex, that is, it parameterizes the lines that are singular points of the intersection of $Y$ with $\alpha$-planes. We can identify each point $x\in X$ with the $\alpha$-plane $\Omega(x)$. An automorphism $g_{i}$ acts on $X$ by sending $\Omega(x)$ to $\Omega(\Pi)$. Let $x\in X$ be a general point and let $\Pi=\mathbb{T}_{x}(X)$ be the embedded tangent space of $X$ at $x$. As we explained above, $g_{i}$ interchanges the two families $\{\Omega(x)\}_{x\in\mathbb{P}^{3}}$, $\{\Omega(\Pi)\}_{\Pi\subset\mathbb{P}^{3}}$ of planes, and hence,
@@ 4. Birational involutions of a quartic surface
where $P=[a,b,c,d],x=[x_{0},x_{1},x_{2},x_{3}],T(x)=[y_{0},y_{1},y_{2},y_{3}]$. It is expected to be a curve of degree $12k+$ 1, where $k$ is the algebraic degree of $T$ [11, p. 175]. However, it also includes the locus of fundamental points of $T$, as well as the closure of the locus of fixed points of $T$. The class of $S(\sigma)$ is expected to be equal to the degree of $\Pi\cap T(\Pi)\cap X$ for a general plane $\Pi$, which is $4k$.
@@ 5. Quartic surfaces: $p=2$
It is obvious that, in characteristic $p\neq 2$, the bitangent surface of a normal quartic surface $X$ does not contain $\alpha$-planes (otherwise, the projection of $X$ from some point is ramified at every point). This is not anymore true over in characteristic $p=2$.
A point $x$ in $\mathbb{P}^{3}$ is called an inseparable projection center of a normal surface $X$ if the projection map with the center at $x$ is inseparable. It is clear that the set of lines passing through an inseparable projection center is an $\alpha$-plane contained in ${\mathrm{Bit}}(X)$. Conversely, if ${\mathrm{Bit}}(X)$ contains an $\alpha$-plane of lines through a point $x\in\mathbb{P}^{3}$, the points $x$ is an inseparable projection center of $X$.
@@
The set of inseparable projection centers of a normal quartic surface in characteristic $2$ is a finite set. Any inseparable projection center contained in $X$ is a singular point of $X$.
@@
Since the set of points $x\in X$ such that the line $\langle q,x\rangle$ is tangent at $x$ is equal to the intersection $X\cap P_{q}(X)$ [3, Theorem 1.1.5], a point $q=[a_{0},a_{1},a_{2},a_{3}]$ is an inseparable center of a quartic surface $X=V(F)$ if and only if the polar $P_{q}(X)=V(\sum a_{i}\frac{\partial F}{\partial x_{i}})=\mathbb{P}^{3}$. Suppose the set of inseparable projection centers contains a curve $Z$ of degree $\geq 2$ (not necessarily irreducible). Then, we can choose three non-collinear points on it. Choose projective coordinates such that these points are $[1,0,0,0],[0,1,0,0]$ and $[0,0,1,0]$. Then, the partials $\frac{\partial F}{\partial x_{i}}=0$ for $i=0,1,2$. This means that $F$ contains $x_{0},x_{1},x_{2}$ in even power. But, then, it must contain $x_{3}$ in even power. Hence, $F$ is a square. So, $Z$ must be a line. In this case, ${\mathrm{Bit}}(X)$ contains a special hyperplane section of the Grassmannian $G_{1}(\mathbb{P}^{3})$ that consists of lines intersecting $Z$. A general plane section of $X$ intersects $Z$ and, hence, contains a curve of bitangent lines. However, a smooth quartic curve contains only finitely many bitangents (in fact, less than or equal to $7$, see below Proposition 5.8). Thus, a general plane section of $X$ is singular contradicting the Akizuli-Bertini theorem (see [9, Remark 8.18.1]). This proves that the set of inseparable centers is a finite set.
@@
We can extend the proof of Proposition 2.2 to the case of characteristic $2$ to obtain the following Corollary.
@@
${\mathrm{Bit}}(X)$ is a congruence of lines, i.e., each irreducible component of ${\mathrm{Bit}}(X)$ is a surface.
Recall that a universal binary form of degree $d$ is a homogeneous polynomial of degree $d$ in two variables whose coefficients are algebraically independent over a field $k$..
@@
By Corollary 2.5, any surface in characteristic zero with ordinary double points as singularities is a general projection surface.
In the next section, we show that it is not true anymore in characteristic $2$.
@@
where $Q$ is a quadratic form and $F$ is a cubic form. The singular locus of $X$ contains $V(L,Q,F)\cup(V(Q)\cap\textrm{Sing}(V(F)))$, where $\textrm{Sing}(V(F))$ is the singular locus of $V(F)$. In particular, $X$ is a singular quartic surface.
@@
Without loss of generality, we may assume that $L=x$ is a coordinate plane. Taking the partials, we find the singular locus contains $V(x,Q,F)\cup(V(Q)\cap\textrm{Sing}(V(F)))$. ∎
Recall that the $p$-rank of a smooth curve $C$ of genus $g>0$ over an algebraically closed field of characteristic $p>0$ is the $p$-rank of the elementary abelian group $J(C)[p]$ of $p$-torsion points of its jacobian variety of $C$. It takes values in the set $[0,g]$. When $p=2$ and $C$ is a smooth plane quartic, the $p$-rank takes values in $\{3,2,1,0\}$ and can be characterized by the number of bitangents of $C$ equal to $7,4,2,1$, respectively [20], [24, §3].
The proof of the following proposition can be found in [28, Proposition 1] (see also [20]).
@@
The loc. cit. paper of Wall also computes the automorphism group of a plane quartic in one of the normal forms (1)-(4). This implies that the codimension of the subspace of plane quartics of the form (1)-(4) is equal to $0,1,2,3$, respectively [28].
Based on Wall’s computations, the following proposition is proved in [19, Section 2].
@@
A smooth plane quartic curve over any field $k$ (not necessary algebraically closed) of characteristic $2$ of $p$-rank less than $3$ is isomorphic over $K$ to a curve $Q^{2}+LF$, where $L$ is a linear form.
The next theorem shows, surprisingly for us, that, although the moduli space of non-ordinary (i.e., of $p$-rank $<3$) is of dimension $5$ in the moduli space of all plane quartics, a general hyperplane section of a nonsingular quartic surface is an ordinary curve.
Theorem 5.10.
Let $X$ be a normal quartic surface over an algebraically closed field of characteristic $2$ whose general hyperplane section is not an ordinary plane quartic. Then, ${\mathrm{Bit}}(X)$ contains a $\beta$-plane, and hence, $X$ is singular.
@@
$F:=\{(x,\Pi)\in X\times\check{\mathbb{P}}^{3}:x\in\Pi\}\subset\{(x,\Pi)\in\mathbb{P}^{3}\times\check{\mathbb{P}}^{3}:x\in\Pi\}$
be the universal family of plane sections of $X$ considered as the closed subset of the universal family of planes. Passing to the generic fiber of the second projection, we obtain a plane quartic curve $C_{K}$ over the field $K$ of rational functions on ˇ$\check{\mathbb{P}}^{3}$. Applying Proposition 5.9, we find that $C_{K}=V(Q^{2}+LF)$, where $L$ is a linear form with coefficients in $K$. Applying a linear transformation of $\mathbb{P}^{3}$ over $K$, we may assume that the equation of $V(L)$ in $\mathbb{P}_{K}^{3}$ is $x_{0}=0$, where $(x_{0},x_{1},x_{2},x_{3})$ are coordinates in $\mathbb{P}_{K}^{3}$. This implies that any line in the plane $V(x_{0})$ is a bitangent line of $X$. Applying Lemma 5.7, we obtain that $X$ is singular. ∎
@@
The surface $X$ has a unique (non-rational) singular point $[0,0,0,1]$ and an ordinary node $[0,0,1,0]$. The point $[0,0,0,1]$ is a unique inseparable projection center of $X$. In fact, take a general point $P=[x_{0},y_{0},z_{0},w_{0}]$ in $\mathbb{P}^{3}$. Substituting the parametric equation $[x,y,z,w]=[sx_{0}+tu_{0},sy_{0}+tu_{1},sz_{0}+tu_{2},sw_{0}+tu_{3}]$, we find one bitangent line that connects $P$ with the point $[0,0,0,1]$. Thus, ${\mathrm{Bit}}(X)$ contains an $\alpha$-plane. On the other hand, a general hyperplane section $w=\alpha x+\beta y+\gamma z$ is the case of Proposition 5.8, that is, it is a plane quartic curve with 2-rank 0. By Proposition 5.10, it contains a $\beta$-plane. Thus, ${\mathrm{Bit}}(X)$ contains the union of an $\alpha$-plane and a $\beta$-plane. Further computation shows that there is nothing else.